# Interlayer exciton condensates between second Landau level orbitals in double bilayer graphene

### v1

@@ Abstract
An interesting question is how to understand the strong $V_{\text{int}}$ dependence of the $N=1$ interlayer EC. We note that the interlayer bias allows us to independently tune $D_{\text{top}}$ and $\nu_{\text{top}}$ in the top BLG while keeping the bottom BLG $\nu_{\text{bot}}$ constant (see SI S6 for details). Fig. 4a and 4b present 2D maps of $R_{\text{xx}}^{\text{drive}}$ and $R_{\text{xx}}^{\text{drag}}$, respectively, for fixed $\nu_{\text{bot}}=$ 3.5, while Fig. 4c shows $R_{\text{xx}}^{\text{drive}}$ for fixed $\nu_{\text{bot}}=$ 1.5. At these two fillings and given our sweeping ranges, we are keeping the bottom BLG mostly in $N=1$ LLs (see SI S6). Fig. 4a and 4c both exhibit resistive patterns that are generally periodic in $\nu_{\text{top}}$, reflecting the QH states of the top BLG that manifest in the bottom BLG $R_{\text{xx}}^{\text{drive}}$ due to screening effects. This is illustrated in Fig. 4d as the four colored stripes bounded by solid lines. More importantly, Fig. 4a and 4c both exhibit four curved boundaries (represented as dashed lines in Fig. 4d) that cut through the repeating LL features near integer $\nu_{\text{top}}$. These curves are consistent with the hierarchy transitions in the top BLG, allowing us to assign each LL section with colors corresponding to its valley flavor and orbital number defined in Fig. 1c.
@@
Among these features, the N = 1 EC states manifest as non-zero drag resistance peaks in the otherwise mostly zero $R_{\text{xx}}^{\text{drag}}$ (Fig. 4b) that align with the resistance dips in $R_{\text{xx}}^{\text{drive}}$ (Fig. 4a and 4c). Note that an EC should exhibit zero resistance in both $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xx}}^{\text{drive}}$. However, an imperfect positioning of $\nu_{\text{bot}}$ can shift $R_{\text{xx}}^{\text{drag}}$ away from the narrow resistance dip and produce the peaks in Fig. 4b. In Fig. 4d, we represent the experimentally observed $N=1$ EC states as thick solid lines. As can be seen immediately, these interlayer EC states occur only when the upper BLG is in the $N=1$ orbital and $K^{\prime}$ valley. We can also demonstrate that, over the same range of gate parameters, the lower BLG remains $N=1$ orbital and $K$ valley (see SI S6). Being an admixture of the conventional lowest and second LL wavefunctions, the $N=1$ orbital with $K^{\prime}$ ($K$) valley in the upper (lower) BLG has its second LL wavefunction component (also has a higher wavefunction weight) polarized to the bottom (top) graphene layer within the BLG (see the upper inset illustration in Fig. 4d). Taken together, our observations suggest that the $N=1$ EC states select a particular flavor polarization so that the higher LL wavefunctions in the two BLGs lie in immediate proximity to each other across the hBN spacer.
@@
Formation of an interlayer EC relies on the interlayer Coulomb interaction and therefore in principle favors bringing the active orbitals as close as possible provided there is no significant interlayer tunneling. However, switching layer polarization of a BLG orbital moves the wavefunction by only 0.34 nm, insignificant compared to the 2.5 nm hBN spacer that sets the interlayer Coulomb interaction scale. Geometric proximity alone is therefore unlikely to explain why the $N=1$ EC selects a specific valley (or layer) flavor combination. A plausible mechanism likely involves the interlayer Haldane pseudopotential, which sensitively depends on the detailed wavefunction distributions. It is possible that the $N=1$ EC is stabilized only in configurations where the second LL component of the $N=1$ LL in one BLG can interact directly with that in the other, free from screening by the co-existing lowest LL component.
@@
Figure 1: a, Device picture with dashed lines denoting the original flake placement. The semi-transparent region is covered by contact gate. b, Schematic of the double layer structure and device configuration. c, Illustration of valley and orbital flavor of the zero-energy Landau levels in BLG. Spin flavor is not included. d, Drive resistance $R_{\text{xx}}^{\text{drive}}$ (bottom layer) as a function of displacement field $D_{\text{bot}}$ and filling factor $\nu_{\text{bot}}$. e, Schematic of the main features in d.
Figure 2: a, $R_{\text{xx}}^{\text{drive}}$ and b, $R_{\text{xx}}^{\text{drag}}$ as a function of $\nu_{\text{top}}$ and $\nu_{\text{bot}}$. c, $R_{\text{xx}}^{\text{drag}}$ near $(\nu_{\text{top}},\nu_{\text{bot}})=(0.5,2.5)$ at $B=16$ T. d, same as c but with $B=25$ T. e, $R_{\text{xx}}^{\text{drag}}$ near $(\nu_{\text{top}},\nu_{\text{bot}})=(0.5,0.5)$ at $B=16$ T. f, Linecut across the EC state in c showing $R_{\text{xx}}^{\text{drag}}$ as well $R_{\text{xy}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drive}}$.
Figure 3: a, $R_{\text{xx}}^{\text{drive}}$ as a function of $\nu_{\text{top}}$ and $\nu_{\text{bot}}$ at $V_{\text{int}}=0.07$ V. Dashed squares label interlayer states formed by N = 0 orbitals. Dashed circles marks the newly observe states formed by $N=1$ orbitals. b, The corresponding drag signal $R_{\text{xx}}^{\text{drag}}$. The top insets show cartoons illustrating the two classes of interlayer states formed by coupling $N=$ 0 and 1 $L$ orbitals. c, $R_{\text{xx}}^{\text{drag}}$ and e, $R_{\text{xy}}^{\text{drag}}$ near $(\nu_{\text{top}},\nu_{\text{bot}})=(1.5,0.5)$ at $V_{\text{int}}=0.1$ V. d and f, Linecuts across the state showing all other transport channels. Dashed lines denote the state location and quantization value.
Figure 4: a, $R_{\text{xx}}^{\text{drive}}$ at fixed $\nu_{\text{bot}}=3.5$ as a function $D_{\text{top}}$ and $\nu_{\text{top}}$. b, $R_{\text{xx}}^{\text{drag}}$ in the same range. Showing clearly interlayer QH states. c, $R_{\text{xx}}^{\text{drive}}$ at fixed $\nu_{\text{bot}}=1.5$. d, Schematic of the features in the resistance data. This shows the $N=1$ interlayer QH states only forms when $N=1$ orbitals are on the nearest layers.
@@ Acknowledgements.
Supplementary Information
@@
Fig. S1a shows the schematic diagram of the device again. $V_{\text{TG}},V_{\text{BG}},V_{\text{int}}$ are voltages applied to the top graphite gate, bottom graphite gate, and top bilayer graphene (BLG), respectively, with the bottom BLG grounded. Taking $V_{\text{est}}$, $V_{\text{esb}}$ and $V_{\text{esint}}$ as electrostatic potential for the top, bottom, and interlayer capacitors, respectively, we have the following equilibrium equalities.
@@
$en_{\text{bot}}=q_{2}-q_{3}=C_{b}(V_{BG}-\mu_{b})-C_{\text{int}}(-V_{\text{int}}+\mu_{b}-\mu_{t})$
@@
$D_{\text{bot}}=\frac{q_{2}+q_{3}}{2}=\frac{1}{2}\left(C_{b}(V_{BG}-\mu_{b})+C_{\text{int}}(-V_{\text{int}}+\mu_{b}-\mu_{t})\right)$
When determining the rough location in the displacement field-density map, we can often first ignore the chemical potential terms. Then the above formulas recover to the ones in the main text.
@@
Figs. S1b and c show the charge neutrality resistance peaks (CNPs) in the two BLG layers, respectively, where we measure their longitudinal resistance $R_{\text{xx}}$ separately using only the top or bottom gate. The sharp peaks demonstrate the high quality of the individual BLG samples. Fig. S2 show transverse and longitudinal resistance $R_{\text{xy}}$ and $R_{\text{xx}}$ in the bottom layer. Clean data with good quantization in $R_{\text{xy}}$ and the corresponding zero $R_{\text{xx}}$ again demonstrate the high quality of the sample.
When converting the gate voltages to $D$ or $n$ associated with the BLG layers, it is important to know the thicknesses of the BN dielectric layers well. Although atomic force microscopy (AFM) can reliably measure relatively thick BN layers, giving the thickness of the top BN (between the top gate and top BLG) thickness $\approx$17 nm and bottom BN (between the bottom gate and bottom BLG) $\approx$12 nm, it can be difficult to accurately measure a thin BN layer. In our experiment, we utilize the principle of measuring the chemical potential of one graphene layer using an adjacent graphene layer to extract the thickness of the BN spacer layer between layers.
@@
$en_{\text{top}}=C_{t}V_{\text{TG}}+\frac{C_{b}(C_{t}+C_{\text{int}})}{C_{\text{int}}}V_{BG}$
$\mu_{t}=-V_{\text{int}}-\frac{C_{b}}{C_{\text{int}}}V_{BG}$
These formulas mean that if we follow the CNP in the bottom BLG — for instance, the resistance peak in Fig. S3a and b — we can directly measure the $n_{\text{top}}$ and $\mu_{t}$ at each point along the peak using the gate voltages at that point and geometric capacitance values. CNP peak in Fig. S3b is shifted to the lower left compared to Fig. S3a due to a finite $V_{BG}$. Using this shift, we can directly extract $C_{b}/C_{\text{int}}$ using the formulas above and obtain an accurate value of the BN spacer layer thickness. Similarly, by tracking the CNP in the top layer, as shown in Fig. S3b and c, we can also extract $C_{t}/C_{\text{int}}$, allowing a consistent check of the relative thicknesses between the BN layers.
III S3. Additional discussion on our measured BLG QH phase diagram
In the main text, we discussed that the presence or absence of the the denominator-3 FQH state features provides a visual distinction can help us roughly identify the orbital number $N=$ 0 or 1 for each region in the $D$ versus $\nu$ phase diagram. However, more accurate identification of the orbital number relies on calculating $D$ and $\nu$, and using the QH phase diagram of single BLG as a reference, from which the flavors of both layers can be deduced.
@@
A key difference between our BLG phase diagram and previous studies is that we did not observe any even-denominator FQH states in N = 1 LLs. Given that the device is encapsulated by both top and bottom graphite gates and fabricated under the cleanest conditions possible, and considering that we observe interlayer FQH states as discussed later, the lack of even-denominator states is less likely due to sample quality but more likely due to the enhanced screening effects from the relatively thin hBN dielectric layers — top hBN is 12 nm ($12~\text{nm}/l_{B}=1.87$ at 16 T, $l_{B}=\sqrt{h/(2\pi eB)}$ is the magnetic length) and bottom hBN is 17 nm. Previous studies showing even-denominator FQH states generally used thicker (>> 35 nm) hBN layers.
@@
Fig. S5 ($V_{\text{int}}=$ 0) and S6 ( $V_{\text{int}}=$ 0.07 V) show complimentary resistance channels to those presented in Fig. 2 and Fig. 3 in the main text. Importantly, Fig. S5c and S6c plot the data as a function of the displacement fields of the top and bottom BLG. By reading off the corresponding ($D_{\text{bot}}$, $D_{\text{top}}$) and ($\nu_{\text{bot}}$, $\nu_{\text{top}}$) values for each interlayer state and consulting the single BLG phase diagram, we can unambiguously identify the flavor and orbital of each BLG underlying the interlayer state.
@@
Fig. S8-S12 are zoomed-in scans of various resistance channels for the $N=0$ and $N=1$ EC states, complementing the zoom-in scans shown in the main text Fig. 2 and 3. In particular, the dashed lines in all the e panels of Fig. S8-S10 illustrate the additional interlayer discussed in the main text. These lines well define linear relations with fractional slopes, corresponding to interlayer fractional states or the so-called semiquantized states.

### v2

@@ Abstract
An interesting question is what drives the strong $V_{\text{int}}$ dependence of the $N=1$ interlayer EC. The interlayer bias lets us tune $D_{\text{top}}$ and $\nu_{\text{top}}$ independently while holding $\nu_{\text{bot}}$ constant (see SM, Sec. S6). Fig. 4a and 4b show 2D maps of $R_{\text{xx}}^{\text{drive}}$ and $R_{\text{xx}}^{\text{drag}}$, respectively, at fixed $\nu_{\text{bot}}=$ 3.5, while Fig. 4c shows $R_{\text{xx}}^{\text{drive}}$ at $\nu_{\text{bot}}=$ 1.5. For these fillings and sweep ranges, the bottom BLG remains mostly in $N=1$ LLs (SM, Sec. S6). In Fig. 4a and 4c, $R_{\text{xx}}^{\text{drive}}$ exhibits resistive patterns that are generally periodic in $\nu_{\text{top}}$, reflecting filling of LLs in the top BLG that imprint onto the bottom BLG $R_{\text{xx}}^{\text{drive}}$ via screening effects. Fig. 4d illustrates the partially-filled top BLG LLs as 4 colored stripes bounded by solid lines. More importantly, Fig. 4a and 4c also show four curved boundaries (dashed lines in Fig. 4d) that cut through the repeating LL features near integer $\nu_{\text{top}}$. These curves are consistent with hierarchy transitions in the top BLG, allowing for assignment of the valley and orbital flavor in each LL segment.
@@
Within this parameter space, the N = 1 EC states appear as straight vertical features: drag resistance peaks in the otherwise mostly zero $R_{\text{xx}}^{\text{drag}}$ (Fig. 4b) that coincide with dips in $R_{\text{xx}}^{\text{drive}}$ (Fig. 4a,c). Note that an EC should exhibit zero resistance in both $R_{\text{xx}}^{\text{drag}}$ and $R_{\text{xx}}^{\text{drive}}$. However, small offsets in $\nu_{\text{bot}}$ can shift $R_{\text{xx}}^{\text{drag}}$ away from the narrow resistance dip and produce the peaks in Fig. 4b. In Fig. 4d, we represent the experimentally observed $N=1$ EC states as thick solid lines. What’s immediately clear is that these interlayer EC states occur only when the upper BLG is in the $N=1$ orbital and $K^{\prime}$ valley. Over the same gate range, the lower BLG remains $N=1$ orbital and $K$ valley (see SM, Sec. S6). Being an admixture of the conventional lowest and second LL wavefunctions, the $N=1$ orbital with $K^{\prime}$ ($K$) valley in the upper (lower) BLG has its second LL wavefunction component (and larger wavefunction weight) polarized toward the bottom (top) graphene layer within that BLG (upper inset in Fig. 4d). Taken together, these results suggest that the $N=1$ EC selects a flavor configuration that places the higher-LL wavefunctions in the two BLGs in immediate proximity across the hBN spacer.
@@
Formation of an interlayer EC relies on the interlayer Coulomb interaction and therefore in principle is favored by bringing the active orbitals as close as possible provided there is no significant interlayer tunneling. However, switching layer polarization of a BLG orbital moves the wavefunction by only 0.34 nm, insignificant compared to the 2.5 nm hBN spacer that sets the interlayer Coulomb interaction scale. Geometric proximity alone is therefore unlikely to explain why the $N=1$ EC selects a specific valley (or layer) flavor combination. A plausible mechanism likely involves the interlayer Haldane pseudopotential, which sensitively depends on the detailed wavefunction distributions. It is possible that the $N=1$ EC is stabilized only in configurations where the second LL component of the $N=1$ LL in one BLG can interact directly with that in the other, free from screening by the co-existing lowest LL component.
@@
Figure S1: a, Device picture with dashed lines denoting the original flake placement. The semi-transparent region is covered by contact gates. b, Schematic of the double layer structure and device configuration. c, Illustration of valley and orbital flavor of the zero-energy Landau levels in BLG. Spin flavor is not included. Valley flavor is tied with layer polarization. d, Drive resistance $R_{\text{xx}}^{\text{drive}}$ (bottom layer) as a function of displacement field $D_{\text{bot}}$ and filling factor $\nu_{\text{bot}}$. e, Schematic of the main features in d.
Figure S2: a, $R_{\text{xx}}^{\text{drive}}$ and b, $R_{\text{xx}}^{\text{drag}}$ as a function of $\nu_{\text{top}}$ and $\nu_{\text{bot}}$. c, $R_{\text{xx}}^{\text{drag}}$ near $(\nu_{\text{top}},\nu_{\text{bot}})=(0.5,2.5)$ at $B=16$ T. The colormap limits are chosen to highlight drag resistance and suppress spurious signals due to contact quality.d, same as c but with $B=25$ T. e, $R_{\text{xx}}^{\text{drag}}$ near $(\nu_{\text{top}},\nu_{\text{bot}})=(0.5,0.5)$ at $B=16$ T. f, Line cuts across the EC state in c showing $R_{\text{xx}}^{\text{drag}}$, $R_{\text{xy}}^{\text{drag}}$ and $R_{\text{xy}}^{\text{drive}}$.
Figure S3: a, $R_{\text{xx}}^{\text{drive}}$ as a function of $\nu_{\text{top}}$ and $\nu_{\text{bot}}$ at $V_{\text{int}}=0.07$ V. Dashed squares mark interlayer states formed by N = 0 orbitals. Dashed circles mark interlayer states formed by $N=1$ orbitals. b, The corresponding drag signal $R_{\text{xx}}^{\text{drag}}$. The top insets show cartoons illustrating the two types of interlayer states formed by coupling $N=$ 0 and 1 LL orbitals. c, $R_{\text{xx}}^{\text{drag}}$ and e, $R_{\text{xy}}^{\text{drag}}$ near $(\nu_{\text{top}},\nu_{\text{bot}})=(1.5,0.5)$ at $V_{\text{int}}=0.1$ V. d and f, Line cuts across the state showing all other transport channels. Dashed lines denote the state location and quantization value.
Figure S4: a, $R_{\text{xx}}^{\text{drive}}$ at fixed $\nu_{\text{bot}}=3.5$ as a function $D_{\text{top}}$ and $\nu_{\text{top}}$. b, $R_{\text{xx}}^{\text{drag}}$ in the same range, showing clearly interlayer QH states. c, $R_{\text{xx}}^{\text{drive}}$ at fixed $\nu_{\text{bot}}=1.5$. d, Schematic of the features in the resistance data. This shows the $N=1$ interlayer QH states form only when the second LL component in the $N=1$ orbitals are in proximate layers.
@@ Acknowledgements.
Z.H. and P.K. conceived the experiment. Z.H. fabricated the devices. Z.H. and A.M.Z. performed the measurements and analysis. K.W and T.T. provided the hBN crystals. P.K. supervised the project. All authors discussed the results and contributed to the manuscript.
Supplementary Materials
@@
Fig. S3a shows the schematic diagram of the device again. $V_{\text{TG}},V_{\text{BG}},V_{\text{int}}$ are voltages applied to the top graphite gate, bottom graphite gate, and top bilayer graphene (BLG), respectively, with the bottom BLG grounded. Taking $V_{\text{est}}$, $V_{\text{esb}}$ and $V_{\text{esint}}$ as electrostatic potentials for the top, bottom, and interlayer capacitors, respectively, we have the following equilibrium equalities.
@@
$en_{\text{bot}}=q_{2}-q_{3}=C_{b}(V_{\text{BG}}-\mu_{b})-C_{\text{int}}(-V_{\text{int}}+\mu_{b}-\mu_{t})$
@@
$D_{\text{bot}}=\frac{q_{2}+q_{3}}{2}=\frac{1}{2}\left(C_{b}(V_{\text{BG}}-\mu_{b})+C_{\text{int}}(-V_{\text{int}}+\mu_{b}-\mu_{t})\right)$
Here, $e$ is the electron charge, $\epsilon_{0}$ is the vacuum permittivity, and $C_{\text{t}}$, $C_{\text{int}}$ and $C_{\text{b}}$ are the geometric capacitances between the top graphite and top BLG, between the top and bottom BLG layers, and between the bottom BLG and bottom graphite, respectively. When determining the rough location in the displacement field-density map, such as those in Fig. 1d and c, Fig. 2a and b, Fig. 3a and b, and Fig. 4a-d, we can often first ignore the chemical potential terms. Then the above formulas reduce to:
$n_{\text{top}}=\frac{1}{e}\left(C_{\text{t}}(V_{\text{TG}}-V_{\text{int}})-C_{\text{int}}V_{\text{int}}\right)$
$n_{\text{bot}}=\frac{1}{e}(C_{\text{b}}V_{\text{BG}}+C_{\text{int}}V_{\text{int}})$
$D_{\text{top}}/\epsilon_{0}=\frac{1}{2}(C_{\text{t}}(V_{\text{TG}}-V_{\text{int}})+C_{\text{int}}V_{\text{int}})$
$D_{\text{bot}}/\epsilon_{0}=\frac{1}{2}(-C_{\text{b}}V_{\text{BG}}+C_{\text{int}}V_{\text{int}})$
Figure S1: Left, schematic of the gating configuration for a double layer graphene device. Right, the corresponding capacitance network circuit model.
However, whenever zooming in to specific interlayer state clusters or when an accurate determination of filling factors is needed, such as those in Fig. 2c-f and Fig. 3c-f, this simplified approach becomes insufficient. In those cases, we instead determine the filling factors by calibrating against experimentally observed features in the data. This procedure follows the established practice in previous studies (see for example, the Supplementary Information of [19]), and we describe the method again in detail below.
The chemical potential contribution to the filling factor is often incorporated through a quantum capacitance, $C_{q}=e^{2}\rho$, where $e$ is electron charge and $\rho$ is the density of states. Then our double layer system can be modeled as a capacitance network circuit shown in Fig. S1 with the 4 open ports corresponding to the top and bottom gates and two BLG layers.
Here $C_{q1}$ and $C_{q2}$ are quantum capacitances of the top and bottom BLG, respectively. When the top or bottom BLG has a high density of states and can screen efficiently, its quantum capacitance becomes much larger than the geometric capacitance. In this limit, the quantum capacitance has negligible impact and the carrier densities are solely determined by geometric capacitances, thus the bottom gate does not affect the filling factor in the top BLG due to the perfect screening of the bottom BLG, and similarly no top gate tuning for the bottom BLG.
Figure S2: Zoom-in resistance map as a function of gates converted to top and bottom layer filling factors.
In contrast, when both BLGs are in partially filled LL, their densities of states–and hence their screening capabilities–are finite. Under these conditions, neither layer can fully screen electric fields, allowing both gates to influence the carrier densities and filling factors of both BLG layers. More specifically, if we define capacitance matrix $C_{ij}=n_{i}/V_{j}$ [40], the effect of the bottom gate voltage on the top BLG carrier density can then be expressed $C_{\text{tB}}=n_{\text{top}}/V_{\text{BG}}=C_{\text{int}}C_{\text{bot}}/(C_{\text{int}}+C_{\text{bot}}+\rho_{\text{bot}})$. If we assume, within the same LL, the screening capability represented by the density of states remains approximately constant, then the gate-density relation can be simplified. This is a reasonable assumption because the energy gaps of 2-component FQH states are mostly below 5K and the variation due to energy gaps within a LL is a small perturbation [19]. Under this assumption, the relation between the carrier densities and top and bottom gates (say, when $V_{\text{top}}$ is fixed) is reduced to a linear transformation: $n_{\text{top}}=C_{\text{top}}V_{\text{TG}}+C_{\text{tB}}V_{\text{BG}}$, $n_{\text{bot}}=C_{\text{bot}}V_{\text{BG}}+C_{\text{bT}}V_{\text{TG}}$. In our data, as illustrated in Fig. S2 for $(\nu_{\text{bot}},\nu_{\text{top}})\sim$(0.5, 2.5), we can calibrate off the individual intralayer 2/3 FQH states so that they show as horizontal or vertical features at constant fillings in $\nu_{\text{bot}}-\nu_{\text{top}}$ maps.
@@
Figs. S3b and c show the charge neutrality resistance peaks (CNPs) in the two BLG layers, respectively, where we measure their longitudinal resistance $R_{\text{xx}}$ separately using only the top or bottom gate. The sharp peaks demonstrate the high quality of the individual BLG samples. Fig. S4 shows transverse and longitudinal resistance $R_{\text{xy}}$ and $R_{\text{xx}}$ in the bottom layer. Clean data with good quantization in $R_{\text{xy}}$ and the corresponding zero $R_{\text{xx}}$ again demonstrate the high quality of the sample.
When converting the gate voltages to $D$ or $n$ associated with the BLG layers, it is important to know the thicknesses of the BN dielectric layers well. Although atomic force microscopy (AFM) can reliably measure relatively thick BN layers, giving the thickness of the top BN (between the top gate and top BLG) $\approx$17 nm and bottom BN (between the bottom gate and bottom BLG) $\approx$12 nm, it can be difficult to accurately measure a thin BN layer. In our experiment, we utilize the principle of measuring the chemical potential of one graphene layer using an adjacent graphene layer to extract the thickness of the BN spacer layer between layers.
@@
$en_{\text{top}}=C_{t}V_{\text{TG}}+\frac{C_{b}(C_{t}+C_{\text{int}})}{C_{\text{int}}}V_{\text{BG}}$
$\mu_{t}=-V_{\text{int}}-\frac{C_{b}}{C_{\text{int}}}V_{\text{BG}}$
These formulas mean that if we follow the CNP in the bottom BLG — for instance, the resistance peak in Fig. S5a and b — we can directly measure the $n_{\text{top}}$ and $\mu_{t}$ at each point along the peak using the gate voltages at that point and geometric capacitance values. CNP peak in Fig. S5b is shifted to the lower left compared to Fig. S5a due to a finite $V_{\text{BG}}$. Using this shift, we can directly extract $C_{b}/C_{\text{int}}$ using the formulas above and obtain an accurate value of the BN spacer layer thickness. Similarly, by tracking the CNP in the top layer, as shown in Fig. S5a and b, we can also extract $C_{t}/C_{\text{int}}$, allowing a consistent check of the relative thicknesses between the BN layers.
The hBN spacer thickness is deliberately chosen to minimize the ratio $d/l_{B}$ to enhance interlayer coupling while simultaneously suppressing interlayer tunneling. Our choice of $d$ = 2.5 nm hBN spacer, corresponding to a $d/l_{B}$   0.4 at B = 16 T, is consistent with the previous studies. For example, Liu et al. [23] employed a 2.5nm thick hBN spacer and J. I. A. Li et al. [19] used a 2.7 nm spacer. In both cases, strong interlayer coupling could be achieved, as evidenced by the plethora of interlayer FQH states reported in these papers.
Figure S6: Leakage current $I_{\text{int}}$ through the hBN spacer as a function of interlayer bias voltage $V_{\text{int}}$
On the other hand, previous tunneling studies have shown that for a hBN spacer with d >> 1.5 nm at small biases ($V_{\text{int}}$  1 V), interlayer tunneling current is negligible. This suppression comes from the large hBN bandgap and lattice orientation mismatch between graphene layers [3, 18, 10]. In our experiments, we further suppress tunneling by intentionally misaligning the two bilayer graphene sheets using their natural straight edges. Owing to these considerations, we observe negligible interlayer current for $V_{\text{int}}<$ 0.2 V (see Fig. S6), which is well above the interlayer bias voltage ranges we explored in the main text.
III S3. Additional discussion on our measured BLG QH phase diagram and its possible connection to the absence of interlayer FQH states in N = 1 orbital
In the main text, we discussed that the presence or absence of the denominator-3 FQH state features provides a visual distinction can help us roughly identify the orbital number $N=$ 0 or 1 for each region in the $D$ versus $\nu$ phase diagram. However, more accurate identification of the orbital number relies on calculating $D$ and $\nu$, and using the QH phase diagram of single BLG as a reference, from which the flavors of both layers can be deduced.
@@
Unlike the N = 0 LL in double mono- or bilayer graphene, in our experiment we don’t observe any interlayer FQH states for N = 1 LL. Here are some considerations and what could be done in the future to explore this direction.
The first factor that stands out is screening from nearby gates. As discussed in the manuscript, the top and bottom hBN dielectric layers have a thickness of 12 nm and 17 nm, respectively. At B = 16 T, the ratio between 12 nm and the magnetic length $l_{B}$ is only 1.88. This relatively small separation from the gates significantly enhances screening, leading to substantial modification of the Coulomb interaction scale $e^{2}/l_{B}$. For example, previous studies have shown that proximity to gates can reduce the energy gaps of FQH states [4], and alter the relative strengths of different Haldane pseudopotentials [30], thereby affecting the stability of certain FQH states, such as the Moore-Read state.
In our experiment, for N = 1 Landau level (LL) orbitals, we do not observe intralayer even-denominator states in an individual BLG, and only very weak signatures of odd-denominator FQH states. The suppression of intralayer FQH remains true even when the other layer is well inside an integer quantum Hall gap, where screening from that layer should be minimal. An example at $V_{\text{int}}=0.1$ V is shown in Fig. S8: the top layer is pinned at top=2 integer quantum Hall state, while the bottom layer sweeps through $0<\nu_{\text{bot}}<2$. The dashed arrows indicate weakly developed 4/3 and 5/3 FQH states, with no observable signatures of the 3/2 state.
Figure S8: top: bottom layer (drive layer) $R_{\text{xx}}^{\text{drive}}$ as a function of $\nu_{\text{bot}}$ and $\nu_{\text{top}}$ at $V_{\text{int}}=0.1$ V, focusing on the contrasting appearances between N = 0 and 1 LL in the bottom layer. bottom: line cuts of $R_{\text{xx}}^{\text{drive}}$ and $R_{\text{xy}}^{\text{drive}}$ at a fixed $\nu_{\text{top}}$. Dashed arrows mark the weak development of 4/3 and 5/3 FQH states in the N = 1 LL.
Meanwhile, the majority of quantum Hall studies on BLG that report well-developed even- and odd- denominator FQH states have much thicker hBN dielectric layers. For example: Ke Huang, Jun Zhu, et al. [13] reports devices with hBN layers all greater than 23 nm; Zibrov, Young, et al., [40] 40-50nm BN for all devices; Assouline, Young, et al., [2] 48.9/37.2 nm; Kumar, Ronen, et al., [17] all around 30–40nm.
This comparison suggests that enhanced gate screening in our device, arising from the relatively thin hBN layers, may be responsible for the absence of intralayer and presumably also interlayer FQH states. Future devices incorporating thicker BN dielectrics will be essential to mitigate screening effects and provide better stability in these correlated states. The second contributing factor could be related to device geometry. In the double layer structure of monolayer graphene, improvements in sample quality, which can be enabled by double graphite gates, were crucial in advancing from observing only interlayer quantum Hall states [25, 21] to resolving more intricate interlayer FQH states [23, 19]. Our device is already encapsulated by top and bottom graphite gates, the sample quality therefore is less likely a limiting factor.
Nevertheless, even for devices of comparable quality, Corbino geometry often exhibits a larger number of interlayer FQH states with significantly improved visibility [19]. This is likely because the Corbino geometry eliminates edge transport, which depends sensitively on equilibrium between contacts and edge modes, and is also susceptible to more disordered physical device edges. By suppressing edge contribution, the Corbino devices provide better sensitivity to bulk incompressible states [36]. Therefore, in our current double layer structure of Bernal bilayer graphene, it is possible that additional more fragile interlayer FQH states are present but remain unresolved due to the limited probing sensitivity. Such states, if they exist, can be accessible using more bulk sensitive geometries or probes like Corbino devices. Another important point we wish to emphasize is that a recent Corbino-geometry study of double monolayer graphene [38] revealed a remarkable level of complexity for the interlayer FQH states, comparable to–if not more–the Jain sequence of intralayer FQH states in a monolayer graphene. This result suggests that there is no intrinsic limitation to realizing interlayer FQH states between N = 1 orbitals in double layer BLG, particularly given the robust observations of intralayer FQH states in single BLG layers.
It’s also possible that N = 1 LL intrinsically supports a different, and potentially weaker, interlayer FQH hierarchy structure, in which case further theory exploration is needed. As discussed in the manuscript, the LL index labels orbitals with different wavefunctions, which directly modify the Coulomb interaction and consequently, the stability and nature of the resulting FQH states. For example, [6] presents a theoretical study of interlayer FQH states between N = 0 orbitals, while the theory in Supplementary Information of [33] examines interlayer integer states involving higher orbits (N >> 1). Comparable theoretical treatments specifically addressing interlayer FQH between N = 1 orbitals are still lacking.
Many of the interlayer fractional states are understood as sharing interlayer vortices in the composite fermion picture. Within this framework, strengthening the interlayer correlation is expected to enhance the stability of interlayer quantum Hall states. This consideration motivates future experiments employing devices with an even thinner hBN spacer to increase interlayer coupling.
@@
Fig. S9 ($V_{\text{int}}=$ 0) and S10 ( $V_{\text{int}}=$ 0.07 V) show complementary resistance channels to those presented in Fig. 2 and Fig. 3 in the main text. Importantly, Fig. S9c and S10c plot the data as a function of the displacement fields of the top and bottom BLG. By reading off the corresponding ($D_{\text{bot}}$, $D_{\text{top}}$) and ($\nu_{\text{bot}}$, $\nu_{\text{top}}$) values for each interlayer state and consulting the single BLG phase diagram, we can unambiguously identify the flavor and orbital of each BLG underlying the interlayer state.
@@
Fig. S12S16 are zoomed-in scans of various resistance channels for the $N=0$ and $N=1$ EC states, complementing the zoom-in scans shown in the main text Fig. 2 and 3. We observe fractional states of both intra- and interlayer nature. For example, the vertical streaks of $R_{\text{xx}}^{\text{drag}}=0$ at $\nu_{\text{bot}}=\frac{2}{3}$ in Fig. 2c and 2e, as well as the horizontal streaks of $R_{\text{xx}}^{\text{drag}}=0$ at $\nu_{\text{top}}=2+\frac{2}{3}$ in Fig. 2c and $\nu_{\text{top}}=\frac{2}{3}$ in Fig. 2e correspond to intra-layer FQH states that developed independently in each BLG. There are also $R_{\text{xx}}^{\text{drag}}$ peaks that follow fractional-slope lines corresponding to interlayer FQH states, including $\nu_{\text{bot}}+\frac{2(\nu_{\text{top}}-2)}{3}=1$ and $\nu_{\text{bot}}+\frac{3(\nu_{\text{top}}-2)}{5}=1$ in Fig. 2c, and $\nu_{\text{bot}}+\frac{2\nu_{\text{top}}}{3}=1$, $\nu_{\text{bot}}+\frac{\nu_{\text{top}}}{3}=1$ and $\frac{\nu_{\text{bot}}}{3}+\nu_{\text{top}}=1$ in Fig. 2e. At higher magnetic fields, additional intra- and interlayer FQH states emerge. For example, compared to 16 T (Fig. 2c), the data at 25 T (Fig. 2d) present additional intralayer FQH states at $\nu_{\text{top/bot}}=\frac{1}{3},\frac{3}{5}$, and interlayer FQH states along $\frac{3\nu_{\text{bot}}}{2}+(\nu_{\text{top}}-2)=1$ and $\nu_{\text{bot}}+\frac{3(\nu_{\text{top}}-2)}{2}=1$. In particular, the dashed lines in all the e panels of Fig. S12-S14 illustrate the additional interlayer states discussed in the main text. These lines well define linear relations with fractional slopes, corresponding to interlayer fractional states or the so-called semiquantized states.